# Safe RLHF-V: Safe Reinforcement Learning from Human Feedback in Multimodal Large Language Models

### v1

@@ 3.3 Moderation Application: Beaver-Guard-V
Due to the inherent vulnerability of MLLMs, the guardrail model ensures safe content generation by filtering out harmful user queries and model responses [4]. We utilize the safety meta labels from the BeaverTails-V, including safety taxonomy labels (i.e., safe or unsafe) and multi-level safety meta tags (i.e., minor, moderate, and severe), to fine-tune the Beaver-Guard-V model for detecting diverse forms of harmful content (see Fig. 3). Experiment results show that the Beaver-Guard-V model accurately classifies specific harmful content categories, enabling targeted filtering strategies tailored to different content types.
To evaluate its effectiveness in identifying toxic content, we assess the Beaver-Guard-V model on both binary and multi-class classification tasks. As shown in Tab. 2, the model achieves a high accuracy rate of 86% in binary classification. Notably, it maintains a low false-positive rate, effectively identifying harmful content while minimizing the erroneous flagging of benign responses. This balance is particularly crucial in real-world applications, where false positives can disrupt user experience and compromise model reliability. In the multi-class classification setting, the model also demonstrates strong performance. Supported by the well-annotated Bevertails-V comprising 20 distinct harmful content categories, the model attains an 84% accuracy rate in classifying various types of harmful content. This fine-grained classification capability allows Beaver-Guard-V to manage a broad spectrum of harmful content with high precision.
As shown in Tab. 3, we further evaluate Beaver-Guard-V’s effectiveness in reducing attack success rates (ASR) under different detection rounds (aka, Filter of N; FoN). The results reveal that increasing moderation rounds consistently lower ASR across different MLLMs. Specifically, at a moderation round of 5, Beaver-Guard-V achieves the lowest ASR across all evaluated models. This model can be used to emphasize the importance of adaptive moderation strategies in maintaining the safety of MLLM interactions.
Figure 3: Moderation process in response to harmful queries.
@@ 4 Method: Safe RLHF-V
4.1 Background
Typically, RLHF consists of three stages:
Supervised Fine-tuning
RLHF begins with a pre-trained model, which is then fine-tuned via supervised learning on a high-quality human instruction dataset designed for downstream tasks. This process results in an chat model $\bm{\theta}_{\mathrm{SFT}}$.
Reward Modeling
This phase involves the collection of preference data $\mathcal{D}$ for training the RM $r_{\mathrm{RM}}(\cdot|\cdot)$. The chat model $\bm{\theta}_{\mathrm{SFT}}$ generates response pairs $({\bm{y}}_{1},{\bm{y}}_{2})$ from given prompts ${\bm{x}}$. Human annotators are then tasked with selecting their preferred response from each pair, denoted as ${\bm{y}}_{w}\succ{\bm{y}}_{l}\mid{\bm{x}}$, where ${\bm{y}}_{w}$ and ${\bm{y}}_{l}$ denote the preferred and preferred answer amongst $({\bm{y}}_{1},{\bm{y}}_{2})$. Finally, the RM $r_{\mathrm{RM}}(\cdot|\cdot)$ can be trained by,
@@ 4.1 Background
Policy Optimization via RL
The final step is to optimize chat model via RL, guided by the reward model $r_{\mathrm{RM}}$. The chat model process that generates responses from the prompts is modeled as a bandit setting, where a reward is obtained from $r_{\mathrm{RM}}$ at the end of each response. The primary objective of RL is to adjust the parameters $\bm{\theta}_{\mathrm{LM}}$ of the LLM so that the expected reward on the training prompt distribution $\mathcal{P}$ is maximized. That is,
@@
Safe reinforcement learning aims to ensure that policy learning adheres to safety constraints typically formulated within the CMDP framework [2].
Constrained Markov Decision Process
is defined by the tuple $\left(\mathcal{S},\mathcal{A},R,\mathbb{P},\mu,\gamma,\mathcal{C}\right)$. Here $\mathcal{S}$ is state space, $\mathcal{A}$ is action space, $\mathbb{P}\left(s^{\prime}\mid s,a\right)$ is probability of state transition from $s$ to $s^{\prime}$ after playing $a$. $r(\cdot):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ and $r\left(s^{\prime}\mid s,a\right)$ denote the reward that the agent observes when state transition from $s$ to $s^{\prime}$ after it plays $a$. $\mu(\cdot)$ : $\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$. A stationary parameterized policy $\pi_{\bm{\theta}}$ is a probability distribution defined on $\mathcal{S}\times\mathcal{A},\pi_{\bm{\theta}}(a\mid s)$ denotes the probability of playing $a$ in state $s$. The goal of RL is to maximize the,
@@
$\begin{split}&\pi_{\star}=\operatorname*{arg\,max}_{\pi_{\bm{\theta}}\in\Pi_{\mathcal{C}}}J\left(\pi_{\bm{\theta}}\right),\\ ~{}~{}\text{s.t.}~{}~{}&\mathbb{E}_{\pi_{\bm{\theta}}}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{i}\left(s_{t},a_{t}\right)\right]\leq b_{i}.\end{split}$ (3)
@@ 4.2 Multimodal Reward and Cost Model
Reward Model-Vision (RM-V)
Utilizing the helpfulness dataset $\mathcal{D}_{R}=\left\{\left({\bm{x}}^{i},{\bm{y}}^{i}_{w},{\bm{y}}^{i}_{l}\right)\right\}_{i=1}^{N}$, we train a parameterized RM $R_{\varphi}({\bm{y}},{\bm{x}})$, where $R_{\varphi}({\bm{y}},{\bm{x}})$ represents a scalar output. This model is trained with the pairwise comparison loss,
@@
Cost Model Vision (CM-V)
Unlike the helpfulness human preference dataset, the harmlessness human preference dataset provides additional information about the harmlessness of a response. To make optimal use of this information for training the CM $C_{\varphi}({\bm{y}},{\bm{x}})$, we amend the original pairwise comparison loss by incorporating classification terms, similar to Safe RLHF [5],
$\mathcal{L}_{C}=$ $-\mathbb{E}_{e\sim\mathcal{D}_{C}}\left[\log\sigma\left(C_{\psi}(y_{w},x)-C_{\psi}(y_{l},x)\right)\right]$
$~{}~{}-\textbf{k}*\mathbb{E}_{(x,y_{w},y_{l},s_{w},s_{l})\sim\mathcal{D}_{C}}\left[\log\sigma\left(s_{w}\cdot C_{\psi}(y_{w},x)\right)\right.$
$~{}~{}\left.+\log\sigma\left(s_{l}\cdot C_{\psi}(y_{l},x)\right)\right],$ (5)
@@
Models Beavertails-V MM-SafetyBench SPA-VL
[…]
Llava-1.5-7B 0.5 0.5 0.5 0.5 0.5 0.5
+RLHF-V 0.110 0.562 0.174 0.687 0.207 0.341
+Llava-RLHF 0.738 0.435 0.665 0.514 0.591 0.446
+DPO (Helpful*) 0.851 0.235 0.697 0.529 0.692 0.491
+DPO (Safety*) 0.441 0.793 0.459 0.766 0.339 0.762
+PPO (Helpful) 0.672 0.528 0.563 0.649 0.547 0.593
+PPO (Safety) 0.253 0.957 0.524 0.717 0.433 0.947
+SafeRLHF-V 0.770 0.579 0.774 0.720 0.613 0.620
Qwen2-VL-7B 0.5 0.5 0.5 0.5 0.5 0.5
+RLHF-V* 0.544 0.452 0.447 0.491 0.370 0.322
+Llava-RLHF 0.647 0.549 0.504 0.521 0.441 0.429
+DPO (Helpful) 0.796 0.630 0.595 0.507 0.585 0.557
+DPO (Safety) 0.599 0.831 0.536 0.799 0.599 0.634
+PPO (Helpful) 0.651 0.506 0.533 0.477 0.566 0.602
+PPO (Safety) 0.357 0.883 0.384 0.761 0.456 0.713
+SafeRLHF-V 0.720 0.689 0.578 0.780 0.626 0.683
LLaVA-1.6-7B 0.5 0.5 0.5 0.5 0.5 0.5
+RLHF-V* 0.065 0.421 0.154 0.398 0.250 0.406
+Llava-RLHF 0.449 0.581 0.594 0.638 0.607 0.541
+DPO (Helpful) 0.737 0.494 0.537 0.579 0.519 0.440
+DPO (Safety) 0.379 0.758 0.431 0.676 0.418 0.543
+PPO (Helpful) 0.610 0.451 0.719 0.631 0.658 0.561
+PPO (Safety) 0.443 0.594 0.554 0.592 0.604 0.698
+SafeRLHF-V 0.630 0.580 0.704 0.705 0.679 0.684
Table 4: Quantitative comparison of Safe RLHF-V and other methods. Across various benchmarks, Safe RLHF-V provides the strongest performance guarantee in balancing helpfulness and harmlessness, significantly outperforming RLHF, which relies solely on helpfulness preferences. The contrastive learning mechanism in DPO enables single-dimensional DPO to achieve exceptionally high helpfulness scores. However, its performance in the safety dimension is considerably lower, underscoring the challenge of evaluating the trade-off between helpfulness and harmlessness when RLHF does not explicitly model safety constraints. * indicates that DPO or RLHF was trained using only single-dimensional partial order preferences from Beavertails-V.
4.3 Safe RLHF-V
During the RL phase, we utilize the Reward Model $R_{\phi}$ to estimate the value of human preference for helpfulness, while the Cost Model $C_{\psi}$ for safety. The MLLM we are training is denoted as $\pi_{\theta}(y|x)$. The following optimization objective is a Safe RL scheme previously outlined in, thereby defined as the objective for our Safe RLHF-V setting:
@@ 4.3 Safe RLHF-V
$\min_{\theta}\max_{\lambda\geq 0}[-\mathcal{J}_{R}(\theta)+\lambda\cdot\mathcal{J}_{C}(\theta)],$ (7)
where $\lambda\geq 0$ serves as the Lagrange multiplier.
Optimizing helpfulness $\mathcal{J}_{R}$ often conflicts with minimizing harm $\mathcal{J}_{C}$. Eq. 7 effectively introduces a penalty term, modulated by $\lambda$, to balance these objectives. By iteratively solving the min-max problem, updating both model parameters $\theta$ and $\lambda$, the framework dynamically adjusts to changes in potential harm, preventing imbalanced optimization.

### v2

@@ 3.3 Moderation Application: Beaver-Guard-V
Due to the inherent vulnerability of MLLMs, the guardrail model filters out harmful user queries and model responses to ensure safety generation [46]. We utilize safety taxonomy labels (i.e., safe or unsafe) and multi-level safety meta tags (i.e., minor, moderate, and severe) from the BeaverTails-V to fine-tune the Beaver-Guard-V model for detecting diverse forms of harmful content (see Fig. 2). Experiment results show that the Beaver-Guard-V model accurately classifies specific harmful content categories, enabling targeted filtering strategies tailored to different content types.
To evaluate its effectiveness in identifying toxic content, we assess the Beaver-Guard-V model on both binary and multi-level meta label settings. As shown in Table 2, the model achieves a high accuracy rate of 78% in binary setting. Notably, it maintains a low false positive rate of 47%, effectively identifying harmful content while minimizing the erroneous flagging of benign responses. This balance is particularly crucial in real-world applications, where false positives can disrupt user experience and compromise model reliability.
In the multi-level meta label setting, the model also demonstrates strong performance. Supported by the well-annotated Bevertails-V comprising 20 distinct harmful content categories and multi-level meta label of QA pairs, the model attains an 85% accuracy rate in detecting harmful content. Notably, Beaver-Guard-V attains state-of-the-art results on several benchmarks. Furthermore, Beaver-Guard-V can also classify both the category of harmful content and its level of severity. This fine-grained classification capability allows Beaver-Guard-V to manage a broad spectrum of harmful content with high precision. As shown in Fig. 2, we further evaluate Beaver-Guard-V’s effectiveness in reducing attack success rates (ASR) under different detection rounds (aka, Filter of N; FoN). The results reveal that increasing moderation rounds consistently lower ASR across different MLLMs. Specifically, at a moderation round of 5, Beaver-Guard-V achieves the lowest ASR across all evaluated models. This model can be used to emphasize the importance of adaptive moderation strategies in maintaining the safety of MLLM interactions.
@@ 4 Method: Safe RLHF-V
4.1 Background and Preliminaries
Supervised Fine-tuning.  RLHF begins with a pre-trained model, which is then fine-tuned via supervised learning on a high-quality human instruction dataset designed for downstream tasks. This process results in an chat model $\bm{\theta}_{\mathrm{SFT}}$.
Reward Modeling.  Then, we use preference data $\mathcal{D}$ to train the RM $r_{\mathrm{RM}}(\cdot|\cdot)$. The chat model $\bm{\theta}_{\mathrm{SFT}}$ generates response pairs $({\bm{y}}_{1},{\bm{y}}_{2})$ from given prompts ${\bm{x}}$. Human annotators are then tasked with selecting their preferred response from each pair, denoted as ${\bm{y}}_{w}\succ{\bm{y}}_{l}\mid{\bm{x}}$, where ${\bm{y}}_{w}$ and ${\bm{y}}_{l}$ denote the preferred and preferred answer amongst $({\bm{y}}_{1},{\bm{y}}_{2})$. So, the RM $r_{\mathrm{RM}}(\cdot|\cdot)$ can be trained by,
@@ 4.1 Background and Preliminaries
RL Fine-tuning.  Finally, we optimize the LLM via RL, guided by the reward model $r_{\mathrm{RM}}$, where a reward is obtained from $r_{\mathrm{RM}}$ at the end of each response. The primary objective of RL is to adjust the LLM parameters $\bm{\theta}_{\mathrm{LM}}$ to maximize the expected reward on the prompt distribution $\mathcal{P}$. That is,
@@
Safe reinforcement learning aims to ensure that policy learning adheres to safety constraints typically formulated within the CMDP framework [47, 48].
Constrained Markov Decision Process is defined by the tuple $\left(\mathcal{S},\mathcal{A},R,\mathbb{P},\mu,\gamma,\mathcal{C}\right)$. Here $\mathcal{S}$ is state space, $\mathcal{A}$ is action space, $\mathbb{P}\left(s^{\prime}\mid s,a\right)$ is probability of state transition from $s$ to $s^{\prime}$ after playing $a$. $r(\cdot):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ and $r\left(s^{\prime}\mid s,a\right)$ denote the reward that the agent observes when state transition from $s$ to $s^{\prime}$ after it plays $a$. $\mu(\cdot)$ : $\mathcal{S}\rightarrow[0,1]$ is the initial state distribution and $\gamma\in(0,1)$. A stationary parameterized policy $\pi_{\bm{\theta}}$ is a probability distribution defined on $\mathcal{S}\times\mathcal{A},\pi_{\bm{\theta}}(a\mid s)$ denotes the probability of playing $a$ in state $s$. The goal of RL is to maximize the,
@@
$\pi_{\star}=\operatorname*{arg\,max}_{\pi_{\bm{\theta}}\in\Pi_{\mathcal{C}}}J\left(\pi_{\bm{\theta}}\right),~{}~{}\text{s.t.}~{}~{}\mathbb{E}_{\pi_{\bm{\theta}}}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{i}\left(s_{t},a_{t}\right)\right]\leq b_{i}.$ (3)
@@ 4.2 Multimodal Reward and Cost Model
Reward Model-Vision (RM-V).  Utilizing the helpfulness dataset $\mathcal{D}_{R}=\left\{\left({\bm{x}}^{i},{\bm{y}}^{i}_{w},{\bm{y}}^{i}_{l}\right)\right\}_{i=1}^{N}$, we train a parameterized RM $R_{\varphi}({\bm{y}},{\bm{x}})$, where $R_{\varphi}({\bm{y}},{\bm{x}})$ represents a scalar output. This model is trained with the pairwise comparison loss,
@@
Cost Model Vision (CM-V).  Unlike the helpfulness human preference dataset, the harmlessness human preference dataset provides additional information about the harmlessness of a response. To make optimal use of this information for training the CM $C_{\varphi}({\bm{y}},{\bm{x}})$, we amend the original pairwise comparison loss by incorporating classification terms, similar to Safe RLHF [16],
$\begin{split}\mathcal{L}_{C}=&-\mathbb{E}_{e\sim\mathcal{D}_{C}}\left[\log\sigma\left(C_{\psi}(y_{w},x)-C_{\psi}(y_{l},x)\right)\right]\\ &~{}~{}-\textbf{k}*\mathbb{E}_{(x,y_{w},y_{l},s_{w},s_{l})\sim\mathcal{D}_{C}}\left[\log\sigma\left(s_{w}\cdot C_{\psi}(y_{w},x)\right)+\log\sigma\left(s_{l}\cdot C_{\psi}(y_{l},x)\right)\right],\end{split}$ (5)
@@
4.3 Reinforcement Learning from Multimodal Human Feedback
During the RL phase, we utilize the Reward Model $R_{\phi}$ to estimate the value of human preference for helpfulness, while the Cost Model $C_{\psi}$ for safety. The following optimization objective is a Safe RL scheme previously outlined in, thereby defined as the objective for our Safe RLHF-V setting:
@@ 4.3 Reinforcement Learning from Multimodal Human Feedback
$\min_{\theta}\max_{\lambda\geq 0}[-\mathcal{J}_{R}(\theta)+\lambda\cdot\mathcal{J}_{C}(\theta)].$ (7)
Unlike in language-only settings, we find that the values of the Lagrange multipliers $\lambda$ significantly affect optimization stability in multimodal scenarios. In practice, naive Lagrange updates can lead to model collapse. To stabilize learning process, we introduce a Budget Bound in the update as follows:
$\lambda\leftarrow\operatorname*{\text{proj}}_{\lambda}[\lambda-\alpha(b-\mathcal{J}_{C}(\theta))],$ (8)
where $\alpha$ is the step size. The projection operator $\operatorname*{\text{proj}}_{\lambda}$ projects $\lambda$ back into the interval $[0,\nu_{\max}]$ where $\nu_{\max}$ is chosen so that $\lambda$ does not become too large. In practice, optimizing helpfulness $\mathcal{J}_{R}$ often conflicts with minimizing harm $\mathcal{J}_{C}$. Eq. 7 introduces a penalty term, modulated by $\lambda$, to balance these objectives. By iteratively solving the min-max problem, updating both model parameters and $\lambda$, the framework dynamically adjusts to changes in potential harm, preventing imbalanced optimization.
Table 3: Comparison of Safe RLHF-V and baselines. Across various benchmarks, Safe RLHF-V provides the strongest performance in balancing helpfulness and safety, significantly outperforming RLHF. We train the single preference algorithm using only single-dimensional preferences. The evaluation of the method is grounded in safety alignment application scenarios. We first consider the safety improvement, followed by the overall enhancement across the two dimensions.
Beavertails-V MM-SafetyBench SPA-VL VLGuard VLSBench
[…]
Llava-1.5-7B N/A 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5
+RLHF-V Single 0.562
+ 0.062
0.110
- 0.390
0.687
+ 0.187
0.174
- 0.326
0.341
- 0.159
0.207
0.462
- 0.038
0.410
0.311
- 0.189
0.142
+Llava-RLHF Single 0.435
- 0.065
0.738
0.514
+ 0.014
0.665
+ 0.165
0.446
- 0.054
0.591
0.487
- 0.013
0.413
0.548 + 0.048 0.645 + 0.145
+DPO (Helpful*) Single 0.235
- 0.265
0.851
0.529
+ 0.029
0.697
+ 0.197
0.491
- 0.009
0.692
0.486
- 0.014
0.667
0.441
- 0.059
0.814
+DPO (Safety*) Single 0.793
+ 0.293
0.441
- 0.059
0.766
+ 0.266
0.459
- 0.041
0.762
+ 0.262
0.339
- 0.161
0.910
+ 0.410
0.239
- 0.261
0.667
+ 0.167
0.482
- 0.018
+PPO (Helpful) Single 0.528 + 0.028 0.672 + 0.172 0.649
+ 0.149
0.563
+ 0.063
0.593 + 0.093 0.547 + 0.047 0.507
+ 0.007
0.562
+ 0.062
0.495
- 0.005
0.672
+PPO (Safety) Single 0.957
+ 0.457
0.253
- 0.247
0.717
+ 0.217
0.524
+ 0.024
0.947
+ 0.447
0.433
- 0.067
0.731 + 0.231 0.585 + 0.085 0.740
+ 0.240
0.453
- 0.047
+MM-RLHF (Helpful) Single 0.382
- 0.118
0.534
0.575 + 0.075 0.660 + 0.160 0.438
- 0.062
0.716
0.555
+ 0.055
0.715
+ 0.215
0.372
- 0.128
0.552
+MM-RLHF (Safety) Single 0.837
+ 0.337
0.345
- 0.155
0.845
+ 0.345
0.421
- 0.079
0.703
+ 0.203
0.352
- 0.148
0.696
+ 0.196
0.369
- 0.131
0.619
+ 0.119
0.348
- 0.152
+SafeRLHF-V Dual 0.579 + 0.079 0.770 + 0.270 0.720 + 0.220 0.774 + 0.274 0.620 + 0.120 0.613 + 0.113 0.852 + 0.352 0.793 + 0.293 0.611 + 0.111 0.814 + 0.314
Qwen2-VL-7B N/A 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5
+RLHF-V* Single 0.452
- 0.048
0.544
0.491
- 0.009
0.447
0.322
- 0.178
0.370
0.477
- 0.023
0.773
0.391
- 0.109
0.761
+Llava-RLHF Single 0.549
+ 0.049
0.647
+ 0.147
0.521
+ 0.021
0.504
+ 0.004
0.429
- 0.071
0.441
0.541
+ 0.041
0.748
+ 0.248
0.455
- 0.045
0.807
+DPO (Helpful) Single 0.630
+ 0.130
0.796
+ 0.296
0.507
+ 0.007
0.595
+ 0.095
0.557
+ 0.057
0.585
+ 0.085
0.377
- 0.123
0.623
0.475
- 0.025
0.859
+DPO (Safety) Single 0.831
+ 0.331
0.599
+ 0.099
0.799 + 0.299 0.536 + 0.036 0.634
+ 0.134
0.599
+ 0.099
0.596
+ 0.096
0.625
+ 0.125
0.530 + 0.030 0.747 + 0.247
+PPO (Helpful) Single 0.506
+ 0.006
0.651
+ 0.151
0.477
- 0.023
0.533
0.602
+ 0.102
0.566
+ 0.066
0.602
+ 0.102
0.612
+ 0.112
0.437
- 0.063
0.859
+PPO (Safety) Single 0.883
+ 0.383
0.357
- 0.143
0.761
+ 0.261
0.384
- 0.116
0.713
+ 0.213
0.456
- 0.044
0.876 + 0.376 0.620 + 0.120 0.523
+ 0.023
0.523
+ 0.023
+MM-RLHF (Helpful) Single 0.458
- 0.042
0.462
0.418
- 0.082
0.522
0.512 + 0.012 0.766 + 0.266 0.380
- 0.120
0.888
0.384
- 0.116
0.826
+MM-RLHF (Safety) Single 0.722
+ 0.222
0.443
- 0.057
0.765
+ 0.265
0.217
- 0.283
0.787
+ 0.287
0.443
- 0.057
0.832
+ 0.332
0.575
+ 0.075
0.685
+ 0.185
0.436
- 0.064
+SafeRLHF-V Dual 0.689 + 0.189 0.720 + 0.220 0.780 + 0.280 0.578 + 0.078 0.683 + 0.183 0.626 + 0.126 0.780 + 0.280 0.873 + 0.373 0.605 + 0.105 0.850 + 0.350
LLaVA-1.6-7B N/A 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5
+RLHF-V* Single 0.421
- 0.079
0.065
0.398
- 0.102
0.154
0.406
- 0.094
0.250
0.721
+ 0.221
0.603
+ 0.103
0.274
- 0.226
0.139
+Llava-RLHF Single 0.581
+ 0.081
0.449
- 0.051
0.638
+ 0.138
0.594
+ 0.094
0.541
+ 0.041
0.607
+ 0.107
0.797 + 0.297 0.648 + 0.148 0.454
- 0.046
0.649
+DPO (Helpful) Single 0.494
- 0.006
0.737
0.579
+ 0.079
0.537
+ 0.037
0.440
- 0.060
0.519
0.461
- 0.039
0.513
0.444
- 0.056
0.698
+DPO (Safety) Single 0.758
+ 0.258
0.379
- 0.121
0.676
+ 0.176
0.431
- 0.069
0.543
+ 0.043
0.418
- 0.082
0.572
+ 0.072
0.503
+ 0.003
0.599
+ 0.099
0.494
- 0.006
+PPO (Helpful) Single 0.451
- 0.049
0.610
0.631 + 0.131 0.719 + 0.219 0.561
+ 0.061
0.658
+ 0.158
0.409
- 0.091
0.625
0.308
- 0.192
0.451
+PPO (Safety) Single 0.594
+ 0.094
0.443
- 0.057
0.592
+ 0.092
0.554
+ 0.054
0.698
+ 0.198
0.604
+ 0.104
0.526
+ 0.026
0.357
- 0.143
0.579
+ 0.079
0.435
- 0.065
+MM-RLHF (Helpful) Single 0.404
- 0.096
0.531
0.518
+ 0.018
0.627
+ 0.127
0.497
- 0.003
0.807
0.497
- 0.003
0.744
0.440
- 0.060
0.717
+MM-RLHF (Safety) Single 0.838
+ 0.338
0.394
- 0.106
0.775
+ 0.275
0.273
- 0.227
0.874
+ 0.374
0.564
+ 0.064
0.721
+ 0.221
0.321
- 0.179
0.722
+ 0.222
0.354
- 0.146
+SafeRLHF-V Dual 0.580 + 0.080 0.630 + 0.130 0.705 + 0.205 0.704 + 0.204 0.684 + 0.184 0.679 + 0.179 0.832 + 0.332 0.722 + 0.222 0.534 + 0.034 0.622 + 0.122